OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 3)(x - 1)(x + 5)?

OpenStudy (anonymous):

solve \[x+3=0\\ x-1=0\\ x+5=0\]

OpenStudy (anonymous):

3,-1,5?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

if \(x+3=0\) then you have \(x=-3\)

OpenStudy (anonymous):

-3,1,-5

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

thank you!!! :)

OpenStudy (anonymous):

yw